# Gamma-ray burst collimated fireballs: long vs short bursts by angular momentum and mass ratio

### v2

@@ 9.1 Intrinsic trends of collimated fireball properties varying with $J/M$
    The macroscopic physics processes of the dynamic interplay between gravitational binding and rotational repelling energy for collimating fireballs is the same for slow- or fast-rotating massive core collapses and binary mergers. The processes and results are characterized by the ratio $J/M$ of total angular momentum $J$ and mass $M$ of systems, which conserve in the entire core collapsing or binary merging process. Such overall ratio $J/M$ possesses the same physical meaning for a rotating core collapsing and binary merging into a rotating core collapsing near the nuclear density when fireballs occur. Therefore, the $J/M$-described macroscopic processes should play a similar key role in fireballs’ collimation for massive core collapses and binary mergers.
@@
    Numerical simulations show binary coalescence has a very different structure from core collapses. In binary coalescence cases, collimated fireballs should form in the most later gravitational collapsing phase of binary coalescence when binary stellar cores merge, becoming a fast-rotating core collapsing around the nuclear density. From these insights, we expect some similarity of fireballs’ collimation between core collapses and binary mergers in terms of their $J/M$. To demonstrate it, full numerical simulations in the GR framework are necessary.
However, these arguments are highly speculative rather than justified. We cannot conclude that the trends apply to fast-rotating collapsing systems and binary mergers relating to GRBs’ observations because of our analytical studies based on a simplified model of a slowly rotating and massive core collapsing in the Newtonian approximation. The key issue is how fireballs collimate in gravitation collapses of fast-rotating systems. It is very complicated to perform analytical studies on collimated fireballs formed in the gravitational collapses of fast-rotating massive core and binary systems in the GR framework. To obtain conclusively the properties of collimated fireballs varying in terms of $J/M$, one has to perform numerical studies in the GR framework to study fast-rotating collapsing systems and binary mergers.
@@ 9.2 Long vs short GRBs and jet angles
    The collimated fireball spectral peak $E_{p}^{j}/E_{p}$ significantly increases up to twice larger than the spherical one, because the photon-pair energy density increases as the collimated fireball volume $V^{j}$ decreases. As a result, the collimated fireball spectral peak $E^{j}_{p}$ increases, total energy $E^{j}_{\gamma}$ and characteristic time scale $R^{j}/c$ decrease as angular momentum $J/M$ increases. These trends might give a possible explanation for two distinct Amati correlations, relating to $E_{p}\propto E_{\gamma}^{\chi}$ (8.6), respectively for long and short bursts’ data scattering on the $E_{p}$-$E_{\gamma}$ plane, see Fig 7 (right) of the reference [47], where two distinct Amati relations are observed for long and short bursts.
@@
    The collimated fireball spectral peak $E^{j}_{p}$ and the peak luminosity $L^{j}_{\gamma}$ increase as angular momentum $J/M$ increases. It implies the Yonetoku correlation, relating to $E_{p}\propto L_{\gamma}^{\chi/(\chi+1)}$ (8.2), is not the same for long and short bursts’ data scattering on the $E_{p}$-$L_{\gamma}$ plane, as shown in Figure 7 (left) of the reference [47].
@@
    The collimatied and sphere fireball sizes’ ratio $R_{\gamma}^{j}/R_{\gamma}\sim{\mathcal{O}}(10^{-1})$. It implies that long bursts duration $T_{90}\sim R_{\gamma}/(2c\Gamma^{2}_{0})$ (8.5) are about 10 times lager than short bursts $T^{j}_{90}\sim R^{j}_{\gamma}/(2c\Gamma_{0j}^{2})$. Here, we assume hydrodynamic expansion ejecta Lorentz factors $\Gamma_{0}/\Gamma_{0j}\sim{\mathcal{O}}(1)$, considering the obtained result $E^{j}_{\gamma}/E_{\gamma}\sim{\mathcal{O}}(1)$ in Table 1 and assuming ejecta mass $M_{\rm ej}$ is the same order of magnitude in spherical and collimated cases, i.e., $M^{j}_{\rm ej}/M_{\rm ej}\sim{\mathcal{O}}(1)$. The ratio $T^{j}_{90}/T_{90}\sim{\mathcal{O}}(10^{-1})$ is qualitatively consistent with two distinct classes of long (soft) and short (hard) bursts’ time-duration data scattering on the $E_{p}$-$T_{90}$ plane, see Figures 1 and 6 (up-right) of the reference [47].
@@
    The trends of numerical results as increasing $J/M$ in Table 1 imply possibly positive or negative correlations between the collimated fireball angle $\theta_{\rm coll}$, burst duration $T_{90}^{j}$, spectral peak $E_{p}^{j}$, total energy $E_{\gamma}^{j}$, luminosity $L_{\gamma}^{j}$. These trends indicate that short GRBs (hard photon spectrum) tend to be emitted from a narrow collimated fireball compared to long GRBs (soft photon spectrum) preferentially from a nearly spherical fireball. In other words, short GRBs from binary coalescences of large angular momenta $J/M$ are more jetting (smaller $\theta_{\rm coll}\sim\theta_{\rm jet}$) than long GRBs from rotating core collapses of small angular momenta. Short GRBs originate from more collimated fireballs with large angular momentum $J_{\gamma}/M$. Long GRBs originate from less collimated fireballs with small angular momentum $J_{\gamma}/M$.
@@
    The above observations on the trends vs angular momentum $J/M$ and Sec. 9.3 discussions on the separatrix $(J/M)_{\rm sepa}$ indicate that short GRB progenitors possess larger $J/M$ values than long GRB progenitors. The separatrix (9.1) of this oversimplified model cannot infer the separatrix $T_{90}\approx 1\sim 2$ seconds between long GRBs (massive core collapses) and short GRBs (binary coalescence). However, following the discussions on the viewpoint (iv) above, we find that $R^{j}_{\gamma}/R_{\gamma}=\tau^{j}_{\gamma}/\tau_{\gamma}\approx 10^{-1}$ corresponds to the line $a^{2}\Omega_{c}^{2}\approx 0.02$ indicated by $\Rightarrow$” in Table 1. It infers collimated fireball angles $\theta_{\rm coll}<5^{\circ}$, thus observed jet angles $\theta^{*}_{\rm jet}<5^{\circ}$ for the cases of short GRBs originating from massive binary coalescence. The separatrix of GRB jet angles should be in short and long GRB data.
@@ 9.3 Angular momentum separatrix of core collapses and binary mergers
In addition, the distinct separation between long and short GRB classes is evident in observations. We attempt to give a possible explanation of this phenomenon consistently with our arguments on the properties of GRBs jet progenitors varying with its $J/M$ in Sec. 9.1. For larger $J/M$ values and rotational repelling energies $U_{c}/|U|$ (5.33), a massive rotating core cannot maintain the equilibrium conditions (5.15) with given material binding energy $\lambda$. Thus, rotating stellar cores split. In this case, the most probable configuration is a binary system. This condition indicates two distinct classes: massive rotating core collapses with small $J/M$ values and binary mergers with large $J/M$ values.
@@
    On the contrary, all materials cannot bind together as a single massive core in collapses when the systems have large $J/M$ or $U_{c}/|U|$ values, and the material binding energy is smaller than the repulsive rotation energy. Therefore, some gravitational binding energy is required to balance rotational energy. Such systems are very probably realised as binary systems undergoing gravitational mergers. Such progenitors produce more collimated fireballs with large angular momenta $J_{\gamma}/M$.
Although we are not able to calculate the separatrix, these discussions give an insight into the distinction between long and short GRBs’ classes relating to the angular momentum and mass ratios $J/M$ of massive rotating core collapses and binary coalescence. The realistic situation of separating two distinct GRB classes should be much more complex. Our theoretical model and analysis methods are too simplified and approximate for fast-rotating core collapses and binary coalescence. Therefore, our physical explanations of jet angle correlation, short GRBs and long GRBs distinction in terms of $J/M$ are speculative. Further theoretical and empirical studies on these explanations are necessary and inviting.

### v3

@@ 9.1 Intrinsic trends of collimated fireball properties varying with $J/M$
    The macroscopic physics processes of the dynamic interplay between gravitational binding and rotational repelling energy for collimating fireballs is the same for slow- or fast-rotating massive core collapses and binary mergers. The processes and results are characterized by the ratio $J/M$ of total angular momentum $J$ and mass $M$ of systems, which are conserved in the entire core collapsing or binary merging process. Such an overall ratio $J/M$ possesses the same physical meaning for a rotating core collapsing and binary merging into a rotating core collapsing near the nuclear density when fireballs occur. Therefore, the $J/M$-described macroscopic processes should play a similar key role in fireballs’ collimation for massive core collapses and binary mergers.
@@
    Numerical simulations show binary coalescence has a very different structure from core collapses. In binary coalescence cases, collimated fireballs should form in the latest gravitational collapsing phase of binary coalescence when binary stellar cores merge, becoming a fast-rotating core collapsing around the nuclear density. From these insights, we expect some similarity of fireballs’ collimation between core collapses and binary mergers in terms of their $J/M$. To demonstrate it, full numerical simulations in the GR framework are necessary.
However, these arguments are highly speculative rather than justified. We cannot conclude that the trends apply to fast-rotating collapsing systems and binary mergers relating to GRBs’ observations because of our analytical studies based on a simplified model of a slowly rotating and massive core collapsing in the Newtonian approximation. The key issue is how fireballs collimate in gravitational collapses of fast-rotating systems. It is very complicated to perform analytical studies on collimated fireballs formed in the gravitational collapses of fast-rotating massive cores and binary systems in the GR framework. To obtain conclusively the properties of collimated fireballs varying in terms of $J/M$, one has to perform numerical studies in the GR framework to study fast-rotating collapsing systems and binary mergers.
@@ 9.2 Long vs short GRBs and jet angles
    The collimated fireball spectral peak $E_{p}^{j}/E_{p}$ significantly increases up to twice as large as the spherical one, because the photon-pair energy density increases as the collimated fireball volume $V^{j}$ decreases. As a result, the collimated fireball spectral peak $E^{j}_{p}$ increases, the total energy $E^{j}_{\gamma}$ and the characteristic time scale $R^{j}/c$ decrease as angular momentum $J/M$ increases. These trends might give a possible explanation for two distinct Amati correlations, relating to $E_{p}\propto E_{\gamma}^{\chi}$ (8.6), respectively for long and short bursts’ data scattering on the $E_{p}$-$E_{\gamma}$ plane, see Fig. 7 (right) of the reference [66], where two distinct Amati relations are observed for long and short bursts.
@@
    The collimated fireball spectral peak $E^{j}_{p}$ and the peak luminosity $L^{j}_{\gamma}$ increase as the angular momentum $J/M$ increases. It implies the Yonetoku correlation, relating to $E_{p}\propto L_{\gamma}^{\chi/(\chi+1)}$ (8.2), is not the same for long and short bursts’ data scattering on the $E_{p}$-$L_{\gamma}$ plane, as shown in Figure 7 (left) of the reference [66].
@@
    The collimated and sphere fireball sizes’ ratio $R_{\gamma}^{j}/R_{\gamma}\sim{\mathcal{O}}(10^{-1})$. It implies that long bursts duration $T_{90}\sim R_{\gamma}/(2c\Gamma^{2}_{0})$ (8.5) are about 10 times lager than short bursts $T^{j}_{90}\sim R^{j}_{\gamma}/(2c\Gamma_{0j}^{2})$. Here, we assume hydrodynamic expansion ejecta Lorentz factors $\Gamma_{0}/\Gamma_{0j}\sim{\mathcal{O}}(1)$, considering the obtained result $E^{j}_{\gamma}/E_{\gamma}\sim{\mathcal{O}}(1)$ in Table 1 and assuming ejecta mass $M_{\rm ej}$ is the same order of magnitude in spherical and collimated cases, i.e., $M^{j}_{\rm ej}/M_{\rm ej}\sim{\mathcal{O}}(1)$. The ratio $T^{j}_{90}/T_{90}\sim{\mathcal{O}}(10^{-1})$ is qualitatively consistent with two distinct classes of long (soft) and short (hard) bursts’ time-duration data scattering on the $E_{p}$-$T_{90}$ plane, see Figures 1 and 6 (up-right) of the reference [66].
@@
    The trends of numerical results as increasing $J/M$ in Table 1 imply possibly positive or negative correlations between the collimated fireball angle $\theta_{\rm coll}$, burst duration $T_{90}^{j}$, spectral peak $E_{p}^{j}$, total energy $E_{\gamma}^{j}$, luminosity $L_{\gamma}^{j}$. These trends indicate that short GRBs (hard photon spectrum) tend to be emitted from a narrow collimated fireball, in comparison with long GRBs (soft photon spectrum) preferentially come from a nearly spherical fireball. In other words, short GRBs from binary coalescence of large angular momenta $J/M$ are more jetting (smaller $\theta_{\rm coll}\sim\theta_{\rm jet}$) than long GRBs from rotating core collapses of small angular momenta. Short GRBs originate from more collimated fireballs with large angular momentum $J_{\gamma}/M$. Long GRBs originate from less collimated fireballs with small angular momentum $J_{\gamma}/M$.
@@
    The above observations on the trends vs angular momentum $J/M$ and Sec. 9.3 discussions on the separatrix $(J/M)_{\rm sepa}$ indicate that short GRB progenitors possess larger $J/M$ values than long GRB progenitors. The separatrix (9.1) of this oversimplified model cannot infer the separatrix $T_{90}\approx 1\sim 2$ seconds between long GRBs (massive core collapses) and short GRBs (binary coalescence). However, following the discussions on the viewpoint (iv) above, we find that $R^{j}_{\gamma}/R_{\gamma}=\tau^{j}_{\gamma}/\tau_{\gamma}\approx 10^{-1}$ corresponds to the line $a^{2}\Omega_{c}^{2}\approx 0.02$ indicated by $\Rightarrow$” in Table 1. It infers collimated fireball angles $\theta_{\rm coll}<5^{\circ}$, thus observed jet angles $\theta^{*}_{\rm jet}<5^{\circ}$ for the cases of short GRBs originating from massive binary coalescence. Analogously to the $T_{90}$ separatrix, the separatrix of GRB jet angles should be in short and long GRB data. This also implies that short and long GRB events should have a bimodal distribution in terms of jet angles  [^5] Thanks to the Referee for this comment..
  7. (vii)
    The above observations are based only on the features of collimated fireballs in the initial prompt era. The GRBs’ ultra-relativistic outflow jet angles $\theta_{\rm jet}\approx\Gamma^{-1}_{\rm jet}\propto(E_{\gamma}/M_{\rm ism})^{-1/8}$ determined by the observed jet breaks in the light curves of the afterglow era [67, 68] for spherical adiabatic evolution, when jets propagate and interact with the surrounding interstellar medium (ISM). The $M_{\rm ism}$ is the total ISM mass engulfed by propagating jets up to the jet breaks’ time $t_{\rm jet}\propto M_{\rm ism}^{1/3}$, and energy-momentum conservation laws are used to determine how the Lorentz factor $\Gamma_{\rm jet}$ decreases from $\Gamma_{0}$ (8.3). We assume that jets conserve the angular momentum along their direction by propagating through a homogeneous ISM without shearing forces. Short GRBs jets from binary coalescence have large $\Gamma_{0}$, small $\theta_{\rm jet}$ and cross-section interacting with ISM, while long GRBs jets from massive core collapses have small $\Gamma_{0}$, large $\theta_{\rm jet}$ and cross-section interacting with ISM. We need to collect more data on jet breaks, particularly for short GRBs events, to see if there is any bimodal distribution in terms of jet angles $\theta_{\rm jet}$, and whether it is consistent with the $T_{90}$ and $E_{p}$ dichotomies for short and long GRBs.
@@ 9.3 Angular momentum separatrix of core collapses and binary mergers
In addition, the distinct separation between long and short GRB classes is evident in observations. We attempt to give a possible explanation of this phenomenon consistently with our arguments on the properties of GRBs’ jet progenitors varying with their $J/M$ in Sec. 9.1. For larger $J/M$ values and rotational repelling energies $U_{c}/|U|$ (5.33), a massive rotating core cannot maintain the equilibrium conditions (5.15) with given material binding energy $\lambda$. Thus, rotating stellar cores split. In this case, the most probable configuration is a binary system. This condition indicates two distinct classes: massive rotating core collapses with small $J/M$ values and binary mergers with large $J/M$ values.
@@
    On the contrary, all materials cannot bind together as a single massive core in collapses when the systems have large $J/M$ or $U_{c}/|U|$ values, and the material binding energy is smaller than the repulsive rotation energy. Therefore, some gravitational binding energy is required to balance rotational energy. Such systems are very probably realized as binary systems undergoing gravitational mergers. Such progenitors produce more collimated fireballs with large angular momenta $J_{\gamma}/M$.
Although we are not able to calculate the separatrix, these discussions give an insight into the distinction between long and short GRB classes relating to the angular momentum and mass ratios $J/M$ of massive rotating core collapses and binary coalescence. The realistic situation of separating two distinct GRB classes should be much more complex. Our theoretical model and analysis methods are too simplified and approximate for fast-rotating core collapses and binary coalescence. Therefore, our physical explanations of jet angle correlation, short GRBs and long GRBs distinction in terms of $J/M$ are qualitative and speculative. Further theoretical and empirical studies on these explanations are necessary and welcome.